OpenStudy (anonymous):

Two times of natural number is subtracted from its square and the result is 35.find the number. A) 5 B) 7 C) 9 D) 6 E) 8 thats the right qst

OpenStudy (anonymous):

\[x^2 - 2x -35 = 0\] \[(x-7)(x + 5)\] x = 7

OpenStudy (anonymous):

x cannot be negative as it is a natural number so -5 is neglected.. x = 7..

OpenStudy (apoorvk):

Let the natural no. be 'x' So it's square = \(x^2\) 2 times the no. = \(2x\) Now, the square of the no. minus twice itself equals 35. So, \(x^2 - 2x = 35\) So, can you solve this quadratic?

OpenStudy (apoorvk):

@shubham.bagrecha one slight error that you made - it would be "x^2 - 2x" and not "2x - x^2".

OpenStudy (anonymous):

Yeah he has done it wrong..

OpenStudy (anonymous):

let the natural no. be x. 2 times natural no . = 2x square of no. = x^2 So, according to the condition x^2 -2x = 35 x^2 - 2x -35 =0 (x-7)(x+5)=0 so, x=7 and x=-5 x=-5 is not posssible as x is a natural no.